# PiRD: Physics-informed Residual Diffusion for Flow Field Reconstruction

### v1

@@ 1 Introduction
As a typical inverse problem, reconstructing spatial fields from sparse sensor data is a highly challenging task, yet its applications are widespread, encompassing numerous fields such as medicine, industrial devices, remote sensing, fluid dynamics, and more [1, 2, 3, 4]. Efforts to improve traditional methodologies for effectively handling sparse input data have to face a balance between computational efficiency and effectiveness [5]. Meanwhile, due to the potential movement and loss of sparse sensors, the fidelity of experimental setups utilizing these collected data is inevitably affected, which leads to structural instability [6]. Moreover, the noise introduced by the sensor measurement process cannot be ignored. All these aspects lead to a complex challenge for reconstruction, which is difficult to meet by traditional methods.
In fluid dynamics, Direct Numerical Simulation (DNS) is a classic method that simulates fluid motion directly by solving the complete Navier-Stokes equations in all scales. To obtain high-fidelity flow field information, DNS requires significant computational resources, especially when solving ill-posed problems or inverse problems related to turbulent flows. Indeed, owing to computational efficiency challenges, several methods employing simplified turbulence models for approximate calculations have been devised, such as Large Eddy Simulation (LES) [7], Reynolds-Averaged Navier-Stokes (RANS) models [8] and hybrid RANS-LES approaches [9]. Nonetheless, these approaches require a trade-off between accuracy and computational expedience. In addition, data assimilation algorithms are proposed to integrate the measurements with the aforementioned numerical models. For example, Wang et al.[10, 11] introduced an adjoint-variational data-assimilation algorithm to reconstruct the flow fields from measurements and enhance traditional methods in the sparse condition. However, these methods also suffer from computational inefficiency, requiring iterative solutions to the Navier-Stokes equations. To mitigate the conflicts between computational cost and simulation accuracy, we aim to draw insights from end-to-end models designed for rapid inference to bridge the gap between simulations and real-world flow field environments, which have demonstrated success in diverse applications such as image processing.
In the field of image processing, the use of machine learning for super-resolution enhancement and data recovery techniques from low-resolution or sparse pixel points is developing rapidly, including algorithms based on convolutional neural networks (CNN) [12, 13], transformer architectures [14, 15, 16], and generative adversarial networks (GAN) [17, 18, 19]. While super-resolution is generally regarded as an image-based data recovery technique, these neural network frameworks have been widely applied to fluid mechanics [20, 21, 22, 23, 24], providing promising alternatives to traditional methods.
@@
Revisiting techniques from the field of image reconstruction, if we regard sparse low-resolution fluid flow data as noisy experimental measurements, then super-resolution flow field reconstruction can be considered as a natural denoising problem. Therefore, the diffusion model in the generation problem [25, 26, 27] also has the potential to reconstruct the sparse data flow field. In this context, Shu et al. [28] introduced a model based on the Diffusion Probabilistic Model (DDPM) for super-resolution tasks, demonstrating its versatility across various scenarios to overcome the limitations in the CNN-based model used in fluid dynamics.
@@
To tackle these challenges, we propose a residual-based diffusion model for flow field super-resolution in this work, inspired by Yue et al.[29] where a diffusion model with residual shifting was presented. Unlike traditional DDPM-based methods, the residual-based diffusion model requires fewer sampling steps and predicts the original high-fidelity flow field directly. This approach enables the practical application of Physics-Informed Neural Networks (PINNs) by integrating the residuals of the governing equations within the objective function, leading to a brand new model called physics-informed residual diffusion (PiRD) model in this paper. Our model not only demonstrates robust performance in reconstructing low-fidelity data across a variety of conditions in the turbulence experiment, but also enhances both training and sampling efficiency—requiring approximately 20 epochs for training and 20 steps for sampling.
Contributions: We make the following key contributions: (i) We present a novel, effective and motivated approach that embeds physical information into the diffusion model to solve the problem of flow field reconstruction under sparse and noisy measurements. (ii) Due to the nature of the ResShift diffusion model, our proposed method can complete the training or inference in about 20 steps, which greatly improves the calculation speed compared with [28], and it is beneficial for future application to actual real-time flow field scenarios. (iii) Through some classical turbulence experiments, PiRD is proved to be more accurate and robust than the CNN-based model for flow field reconstruction on sparse data. Moreover, PiRD can handle various sensor placements and noisy measurements which are not considered in the training data.
@@ 2.1 Problem Setup
The primary objective of this research is to address the limitations of current CNN-based methods in reconstructing high-fidelity (HF) flow fields, denoted as $\omega\sim U_{\text{HF}}$, from low-fidelity (LF) observations, denoted as $\tilde{\omega}\sim U_{\text{LF}}$. Conventional CNN-based models like UNet, represented as $f_{\theta}$, have shown proficiency in dealing with in-distribution low-fidelity data. This process can be described as:
@@
where $\mathcal{D}$ represents certain the down-sampling operation, or a certain method to degrade the high-fidelity flow field, the neural network is essentially trying to learn this operation to reserve this process, and in-distribution data means the high-fidelity data that was corrupted by this down-sampling operation $\mathcal{D}$, or in which operations that are highly similar to it. However, the learned model can only be applicable for the scenarios that are degraded by this specific learned $\mathcal{D}$, if the degraded flow field is not corrupted by it, the reconstruction process still trying to reverse $\mathcal{D}$ which ultimately make the prediction inaccurate. Therefore, to address this problem, we build up a Markov chain between high-fidelity flow fields $U_{\text{HF}}$ and low-fidelity flow fields $U_{\text{LF}}$ that are independent of $\mathcal{D}$, such that we can reconstruct low-fidelity flow fields given any possible $\mathcal{D}$. In addition, physics-informed loss function can also be integrated into such a process to act as not only a physical constraint to make sure that the prediction follows the underlying physical law, but also a guide for the model to infer the information lost in the low-fidelity flow field and ensure robust performance across diverse and challenging scenarios.
@@ 2.2 Physics-informed Residual Diffusion
This section introduces our proposed model used to map the low-fidelity distribution to the high-fidelity distribution through a learned Markov chain and to ensure that the entire process follows the underlying physical law such that we can eventually produce a high-fidelity prediction that also obeys the physical law. In the following, we present the architecture of Physics-informed Residual Diffusion (PiRD) which is composed of two essential modules: (i) Residual Shifting Diffusion Model that could build a Markov chain between LF distributions and HF distributions; (ii) Physics-informed neural networks(PINNs) for fluid mechanics that ensure the predictions obey the underlying physical law.
@@ 2.2.1 Residual Shifting Diffusion Model for Vorticity Field
Residual Shifting Diffusion Model was first proposed by Yue et al. [29] developed for the image super-resolution task. This method diverges from traditional approaches that transition high-fidelity images to Gaussian noise. Instead, it constructs a Markov chain between high-resolution images and low-resolution images, facilitated by adjusting the residual between them. We apply this method as the backbone of our task, expand this ideal and further develop this idea into the application of low-fidelity flow field super-resolution and sparse flow field reconstruction. Let $\omega$ denoted as the high-fidelity flow field, ad $\tilde{\omega}$ as its corresponding low-fidelity flow field, and let $e=\tilde{\omega}-\omega$ denoted the residual between them, this residual does not only contains the spatially related difference between them but also contains the physical related information lost between them. The forward diffusion Markov chain can be described as:
@@
$q(\omega_{t}|\omega,\tilde{\omega})$ $=\mathcal{N}(\omega_{t};\omega+\eta_{t}e,\kappa^{2}\eta_{t}\{I),\quad t=1,2,\ldots,T,$ (2)
@@
$\min_{\theta}\sum_{t}\lVert f_{\theta}(\omega_{t},t)-\omega\rVert^{2}_{2}.$ (6)
@@ 2.2.2 Physics-informed neural networks (PINNs) for fluid mechanics
Within the realm of fluid mechanics, the application of Physics-Informed Neural Networks (PINNs) to the vorticity transport equations exemplifies a sophisticated methodology for solving parametrized partial differential equations (PDEs) that describe complex fluid behaviours. The vorticity transport equation, pivotal for understanding fluid dynamics, models the evolution of vorticity $\omega$, a vector field that characterizes the local spinning motion of the fluid. This equation is succinctly represented as
$\frac{\partial\omega}{\partial t}+\nabla\cdot(\mathbf{u}\omega)=\nu\nabla^{2}\omega,$ (7)
where $\mathbf{u}$ denotes the velocity field, $\nu$ the kinematic viscosity.
This equation embodies the conservation of vorticity in the absence of external influences and lays the foundation for the application of PINNs in fluid mechanics.
The PINNs framework for solving the vorticity transport equation integrates the physics of the problem directly into the learning process of the neural network. By approximating the solution $\hat{\omega}(\mathbf{x},t)$, the network is trained to comply with the governing PDE across the domain $\Omega$ and throughout the time interval $[0,T]$. This training process is facilitated by the formulation of a composite loss function $L$, which encapsulates the adherence of the network’s output to the underlying physical laws and empirical data. The loss function is articulated as
@@
where each term represents a different aspect of the learning criteria: $L_{\text{PDE}}$ penalizes deviations from the vorticity transport PDE, $L_{\text{data}}$ quantifies the discrepancies between the predictions and actual flow measurements, $L_{\text{IC}}$ and $L_{\text{BC}}$ ensure the model’s output matches the initial and boundary conditions, respectively, $\beta_{i}$ denotes the weighting coefficient for each term. To integrate the initial and boundary conditions crucial for accurately modeling fluid behaviour, PINNs incorporate these conditions directly into the loss function. Initial conditions set the stage for the fluid’s state at the beginning of the observation period, whereas boundary conditions dictate the fluid’s behaviour at the domain’s edges, influencing the evolution of the flow over time. The precise specification of these conditions is vital for capturing the physical scenario under investigation, such as the generation and dynamics of vorticity in flows past obstacles.
@@
Figure 1: The architecture of PiRD, during the forward process, a time-step $t$ was randomly drawn from $(0,T]$, then a UNet was used for predicting $\omega$ from $\omega_{t}$ with the objective function that both penalize the element-wise difference but also the physics loss. During the sampling process, the LF field $\tilde{\omega}$ was first mapped to $\omega_{T}$ then followed by the reverse Markov chain and finally predicted HF field $\omega$.
@@ 2.2.3 Physics-informed Residual Diffusion
The traditional Residual Diffusion Model (RDM) operates by constructing a Markov chain, characterizing a residual, $e$, that bridges the gap between the high-fidelity vorticity field, $\omega$, and its low-fidelity counterpart, $\tilde{\omega}$. This residual is dissected into two principal components: $e_{\textit{fidelity}}$, which represents the element-wise discrepancies, and $e_{\textit{physics}}$, which captures deviations from established physical laws. Consequently, the transformation process from $\tilde{\omega}$ to $\omega$ is encapsulated by the equation:
$\omega=\tilde{\omega}+e_{\textit{fidelity}}+e_{\textit{physics}}.$ (9)
@@
Figure 1 delineates the core architecture of the proposed PiRD model. It utilizes a Markov chain methodology to facilitate the reconstruction of high-fidelity flow fields from their low-fidelity counterparts, ensuring adherence to physical laws throughout the reconstruction process.
The loss function employed in PiRD comprises two components: the element-wise loss and the PDE loss. The element-wise is the L2 norm between the prediction $\hat{\omega}$ and the true flow field $\omega$, normalized by the L2 norm of $\omega$, emphasizing the accuracy of individual elements. The PDE loss ensures that the predicted field complies with the governing physical equations, maintaining physical plausibility. Mathematically, the loss function is expressed as:
$\text{Loss}=\alpha\sum_{t}\|f_{\theta}(\omega_{t},\tilde{\omega})-\omega\|_{2}^{2}+\beta\sum_{t}\|G(f_{\theta}(\omega_{t},\tilde{\omega}))\|_{2}^{2},$ (10)
where $\alpha$ and $\beta$ are the weighting parameters that balance the significance of data fidelity and physical consistency, respectively. The differential operator $G$, employed in the loss function, succinctly represents the dynamics captured by the non-dimensional two-dimensional (2D) vorticity transport equation for incompressible fluid flow. This crucial equation is delineated as:
$\frac{\partial\omega}{\partial t}+(u\cdot\nabla)\omega=\frac{1}{Re}\nabla^{2}\omega,$ (11)
where $\omega$ is the vorticity field, $u$ denotes the velocity field, and $Re$ is the Reynolds number, characterizing the flow’s inertial forces relative to its viscous forces. The term $\frac{1}{Re}\nabla^{2}\omega$ accounts for the diffusion of vorticity, with its magnitude scaled inversely by $Re$.
@@
$G(\hat{\omega})=\frac{\partial\hat{\omega}}{\partial t}+(u\cdot\nabla)\hat{\omega}-\frac{1}{Re}\nabla^{2}\hat{\omega}.$ (12)
The ideal outcome, $G(\hat{\omega})=0$, indicates that the predicted vorticity field adheres to the dynamics governed by the vorticity transport equation. Hence, the PDE loss is defined as:
PDE $\text{PDE loss}=\frac{1}{n}\sum_{i=1}^{n}\|G(\hat{\omega}_{i})\|_{2}^{2},$ (13)
which penalizes predictions that deviate from expected fluid behavior, thereby enforcing compliance with fluid dynamics principles.
To effectively compute the time derivative of vorticity, the model utilizes three channels of input corresponding to vorticity fields at consecutive time frames, $t-1$, $t$, and $t+1$, notice that the $t$ here is different with the $t$ in the diffusion model, instead of representing certain time step during the forward or backward process, the $t$ here means $t$ time frame within the whole time span of the dataset. This tri-channel approach enables the approximation of $\frac{\partial\omega}{\partial t}$ via finite difference methods.
@@
In the sampling process, the reverse Markov chain recovers the high-fidelity flow field starting from the low-fidelity observations $\tilde{\omega}$. This process inverts the forward diffusion, iteratively reconstructing the previous time step $\omega_{t-1}$ from the noised version $\omega_{t}$, until the original flow field $\omega_{0}$ is retrieved:
$\omega_{t-1}=\frac{\eta_{t-1}}{\eta_{t}}\omega_{t}+\frac{\alpha_{t}}{\eta_{t}}\omega+\kappa\sqrt{\frac{\eta_{t-1}}{\eta_{t}}\alpha_{t}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,I)$ (15)
We choose UNet as the underlying neural network architecture because of its exceptional performance in capturing complex spatial hierarchies inherent in image-like data, which is analogous to flow fields. With its encoding and decoding pathways, the architecture’s design is well-suited to preserving high-resolution details crucial for reconstructing the intricate structures found in vorticity fields, rendering it an optimal choice for this superresolution task.

### v2

@@ 1 Introduction
As a typical inverse problem, reconstructing spatial fields from sparse sensor data is a highly challenging task, yet its applications are widespread, encompassing numerous fields such as medicine, industrial devices, remote sensing, fluid dynamics, and more [1, 2, 3, 4, 5, 6]. Efforts to improve traditional methodologies for effectively handling sparse input data have to face a balance between computational efficiency and effectiveness [7]. Meanwhile, due to the potential movement and loss of sparse sensors, the fidelity of experimental setups utilizing these collected data is inevitably affected, which leads to structural instability [8]. Moreover, the noise introduced by the sensor measurement process cannot be ignored. All these aspects lead to a complex challenge for reconstruction, which is difficult to meet by traditional methods.
In fluid dynamics, Direct Numerical Simulation (DNS) [9] is a classic method that simulates fluid motion directly by solving the complete Navier-Stokes equations in all scales. To obtain high-fidelity flow field information, DNS requires significant computational resources, especially when solving ill-posed problems or inverse problems related to turbulent flows. Indeed, owing to computational efficiency challenges, several methods employing simplified turbulence models for approximate calculations have been devised, such as Large Eddy Simulation (LES) [10], Reynolds-Averaged Navier-Stokes (RANS) models [11] and hybrid RANS-LES approaches [12]. Nonetheless, these approaches require a trade-off between accuracy and computational expedience. In addition, data assimilation algorithms are proposed to integrate the measurements with the aforementioned numerical models. For example, Wang et al.[13] introduced an adjoint-variational data-assimilation algorithm to reconstruct the flow fields from measurements and enhance traditional methods in the sparse condition. However, these methods also suffer from computational inefficiency, requiring iterative solutions to the Navier-Stokes equations. To mitigate the conflicts between computational cost and simulation accuracy, it is worth investigating the end-to-end learning-based models designed for rapid inference to bridge the gap between simulations and real-world flow field environments, which have demonstrated success in diverse applications such as image processing.
In the field of image processing, the use of machine learning for super-resolution enhancement and data recovery techniques from low-resolution or sparse pixel points is developing rapidly, including algorithms based on convolutional neural networks (CNN) [14, 15, 16], transformer architectures [17, 18, 19], and generative adversarial networks (GAN) [20, 21, 22]. While super-resolution is generally regarded as an image-based data recovery technique, these neural network frameworks have been widely applied to fluid mechanics [23, 24, 25, 26, 27, 28], providing promising alternatives to traditional methods.
@@
Revisiting techniques from the field of image reconstruction, if we regard sparse low-resolution fluid flow data as noisy experimental measurements, then super-resolution flow field reconstruction can be considered as a natural denoising problem. Therefore, the diffusion model in the generation problem [29, 30, 31, 32, 33] also has the potential to reconstruct the sparse data flow field. In this context, Shu et al. [34] introduced a model based on the Diffusion Probabilistic Model (DDPM) [32] for super-resolution tasks, demonstrating its versatility across various scenarios to overcome the limitations in the CNN-based model used in fluid dynamics.
@@
To tackle these challenges, we propose a residual-based diffusion model for flow field super-resolution in this work, inspired by Yue et al.[35] where a diffusion model with residual shifting was presented. Unlike traditional DDPM-based methods, the residual-based diffusion model requires fewer sampling steps and predicts the original high-fidelity flow field directly. This approach enables the practical application of Physics-Informed Neural Networks (PINNs) [36, 37, 38, 39, 40] by integrating the residuals of the governing equations within the objective function, leading to a brand new model called physics-informed residual diffusion (PiRD) model in this paper. Our model not only demonstrates robust performance in reconstructing low-fidelity data across a variety of conditions in the turbulence experiment, but also enhances both training and sampling efficiency—requiring approximately 20 epochs for training and 20 steps for sampling.
We make the following key contributions: (i) We present a novel, effective and motivated approach that embeds physical information into the diffusion model to solve the problem of flow field reconstruction under sparse and noisy measurements. (ii) Due to the nature of the ResShift diffusion model, our proposed method can complete the training or inference in about 20 steps, which greatly improves the calculation speed compared with [34], and it is beneficial for future application to actual real-time flow field scenarios. (iii) Through some classical turbulence experiments, PiRD is proved to be more accurate and robust than the CNN-based model for flow field reconstruction on sparse data. Moreover, PiRD can handle various sensor placements and noisy measurements which are not considered in the training data.
@@ 2.1 Problem Setup
The primary objective of this research is to address the limitations of current CNN-based methods in reconstructing high-fidelity (HF) flow fields, denoted as $\omega\sim U_{\text{HF}}$, from low-fidelity (LF) observations, denoted as $\tilde{\omega}\sim U_{\text{LF}}$. CNN-based models like UNet, represented as $f_{\theta}$, have shown proficiency in dealing with in-distribution low-fidelity data. This process can be described as:
@@
where $\mathcal{D}$ represents certain the down-sampling operation, or a certain method to degrade the high-fidelity flow field, the neural network is essentially trying to learn this operation to reserve this process, and in-distribution data means the high-fidelity data that was corrupted by this down-sampling operation $\mathcal{D}$, or in which operations that are highly similar to it. However, the learned model can only be applicable for the scenarios that are degraded by this specific learned $\mathcal{D}$, as a result, if the degraded flow field is not corrupted by it, the reconstruction process still trying to reverse $\mathcal{D}$ which ultimately make the prediction inaccurate. Therefore, to address this problem, we build up a Markov chain between high-fidelity flow fields $U_{\text{HF}}$ and low-fidelity flow fields $U_{\text{LF}}$ that are independent of $\mathcal{D}$, such that we can reconstruct low-fidelity flow fields given any possible $\mathcal{D}$. In addition, physics-informed loss function can also be integrated into such a process to act as not only a physical constraint to make sure that the prediction follows the underlying physical law, but also a guide for the model to infer the information lost in the low-fidelity flow field and ensure robust performance across diverse and challenging scenarios.
@@ 2.2 Physics-informed Residual Diffusion
This section introduces our proposed model used to map the low-fidelity distribution to the high-fidelity distribution through a learned Markov chain and to ensure that the entire process follows the underlying physical law such that we can eventually produce a high-fidelity prediction that also obeys the underlying governing equation. In the following, we present the architecture of Physics-informed Residual Diffusion (PiRD) which is composed of two essential modules: (i) Residual Shifting Diffusion Model that could build a Markov chain between LF distributions and HF distributions; (ii) Physics-informed neural networks(PINNs) for fluid mechanics that ensure the predictions obey the underlying governing equation.
@@ 2.2.1 Residual Shifting Diffusion Model for Vorticity Field
Residual Shifting Diffusion Model was first proposed by Yue et al. [35] developed for the image super-resolution task. This method diverges from traditional diffusion approaches that transit high-fidelity images to Gaussian noise. Instead, it constructs a Markov chain between high-resolution images and low-resolution images by shifting the residual between them. We apply this method as the backbone of our task, expand this ideal and further develop it into the application of low-fidelity flow field super-resolution and sparse flow field reconstruction. Let $\omega$ denoted as the high-fidelity flow field, ad $\tilde{\omega}$ as its corresponding low-fidelity flow field, and let $e=\tilde{\omega}-\omega$ denoted the residual between them, this residual does not only contains the spatially related difference between them but also contains the physical related information lost between them. The forward diffusion Markov chain can be described as:
@@
$q(\omega_{t}|\omega,\tilde{\omega})$ $=\mathcal{N}(\omega_{t};\omega+\eta_{t}e,\kappa^{2}\eta_{t}I),\quad t=1,2,\ldots,T,$ (2)
@@
$\min_{\theta}\sum_{t}\lVert f_{\theta}(\omega_{t},\tilde{\omega},t)-\omega\rVert^{2}_{2}.$ (6)
For a comprehensive derivation of forward diffusion process and backward inference process, and the design of the noise schedule, please see Section A of the Appendix.
@@ 2.2.2 Physics-informed neural networks (PINNs) for fluid mechanics
Within the realm of fluid mechanics, the application of Physics-Informed Neural Networks (PINNs) to the vorticity transport equations exemplifies a sophisticated methodology for solving parametrized partial differential equations (PDEs) that describe complex fluid behaviours. The Vorticity Transport Equation in fluid dynamics is pivotal in describing the evolution of the vorticity vector $\omega$, which represents the curl of the fluid velocity $u$. This equation is crucial for understanding the rotational behaviors within fluid flows. In applications where comparative analysis across different flow conditions is essential, the non-dimensional form of the vorticity transport equation, incorporating the Reynolds number $Re$, is frequently used. The Reynolds number characterizes the ratio of inertial forces to viscous forces within the fluid, providing insight into the effects of fluid viscosity across various flow regimes. The non-dimensional vorticity transport equation is expressed as:
$\frac{\partial\omega}{\partial t}+(u\cdot\nabla)\omega=\frac{1}{Re}\nabla^{2}\omega$ (7)
In this form, the equation highlights three major phenomena: (1) The Temporal Change $\frac{\partial\omega}{\partial t}$ describes the rate of change of vorticity over time. (2) The Advection term $(u\cdot\nabla)\omega$ represents the transport of vorticity by the velocity field. (3) The Diffusion $\frac{1}{Re}\nabla^{2}\omega$ accounts for the diffusion of vorticity, moderated by the fluid’s viscosity, which is inversely scaled by the Reynolds number. The vorticity transport equation describes the dynamics of vorticity, accounting for its transport and diffusion in fluid flows. This equation sets the stage for employing Physics-Informed Neural Networks (PINNs) in fluid mechanics, leveraging these fundamental principles to enhance model predictions and simulations. The PINNs framework for solving the vorticity transport equation integrates the physics of the problem directly into the learning process of the neural network. By approximating the solution, the network is trained to comply with the governing PDE across the domain $\Omega$ and throughout the time interval $[0,T]$. This training process is facilitated by the formulation of a composite loss function $L$, which encapsulates the adherence of the network’s output to the underlying physical laws and empirical data. The loss function is articulated as
@@
where each term represents a different aspect of the learning criteria: $L_{\text{PDE}}$ penalizes deviations from the vorticity transport PDE, $L_{\text{data}}$ quantifies the discrepancies between the predictions and actual flow measurements, $L_{\text{IC}}$ and $L_{\text{BC}}$ ensure the model’s output matches the initial and boundary conditions, respectively, $\beta_{i}$ denotes the weighting coefficient for each term. To accurately model fluid behavior, PINNs incorporate initial and boundary conditions directly into the loss function. Initial conditions set the stage for the fluid’s state at the beginning of the observation period, whereas boundary conditions dictate the fluid’s behaviour at the domain’s edges, influencing the evolution of the flow over time. The precise specification of these conditions is vital for capturing the physical scenario under investigation, such as the generation and dynamics of vorticity in flows past obstacles.
@@
Figure 1: The architecture of PiRD. During the forward process, a time-step $t$ is randomly drawn from $(0,T]$, then a UNet is used for predicting $\omega$ from $\omega_{t}$ with the objective function that both penalize the element-wise difference but also the physics loss. During the inference process, the LF field $\tilde{\omega}$ is first mapped to $\omega_{T}$ then followed by the reverse Markov chain and finally predicted HF field $\omega$.
@@ 2.2.3 Physics-informed Residual Diffusion
The vanilla Residual Diffusion Model (RDM) operates by constructing a Markov chain, characterizing a residual, $e$, that quantifies the discrepancy between the high-fidelity vorticity field, $\omega$, and its low-fidelity counterpart, $\tilde{\omega}$. This residual is dissected into two principal components: $e_{\textit{fidelity}}$, which represents the element-wise discrepancies, and $e_{\textit{physics}}$, which captures deviations from established physical laws. Consequently, the relationship between $\tilde{\omega}$ and $\omega$ can be represented by the following equation:
$\omega=\tilde{\omega}+e=\tilde{\omega}+(e_{\textit{fidelity}}+e_{\textit{physics}}).$ (9)
@@
Figure 1 illustrates the core architecture of the proposed PiRD model. It utilizes a Markov chain methodology to facilitate the reconstruction of high-fidelity flow fields from their low-fidelity counterparts, ensuring adherence to physical laws throughout the reconstruction process.
The loss function employed in PiRD comprises two components: the element-wise loss and the PDE loss. The element-wise is the L2 norm between the prediction $\hat{\omega}$ and the true flow field $\omega$, normalized by the L2 norm of $\omega$, emphasizing the accuracy of individual elements. The PDE loss ensures that the predicted field complies with the underlying governing equations, to maintain the physical plausibility. Mathematically, the loss function is expressed as:
$\text{Loss}=\alpha\sum_{t}\|f_{\theta}(\omega_{t},\tilde{\omega},t)-\omega\|_{2}^{2}+\beta\sum_{t}\|G(f_{\theta}(\omega_{t},\tilde{\omega},t))\|_{2}^{2},$ (10)
where $\alpha$ and $\beta$ are the weighting parameters that balance the significance of data fidelity and physical consistency, respectively. The differential operator $G$, employed in the loss function, represents the dynamics captured by the non-dimensional two-dimensional (2D) vorticity transport equation for incompressible fluid flow as shown in Equation 7.
@@
$G(f_{\theta}(\omega_{t},\tilde{\omega},t))=\frac{\partial f_{\theta}(\omega_{t},\tilde{\omega},t)}{\partial t}+(u\cdot\nabla)f_{\theta}(\omega_{t},\tilde{\omega},t)-\frac{1}{Re}\nabla^{2}f_{\theta}(\omega_{t},\tilde{\omega},t).$ (11)
The ideal outcome, $G(f_{\theta}(\omega_{t},\tilde{\omega},t))=0$, indicates that the predicted vorticity field adheres to the dynamics governed by the vorticity transport equation. Hence, the PDE loss component is defined as the right hand side of the of Equation 10, which penalizes predictions $f_{\theta}(\omega_{t},\tilde{\omega},t)$ that deviate from its governing equation, thereby enforcing compliance with fluid dynamics principles.
To effectively compute the rate of change of vorticity over time, the model utilizes three channels of input corresponding to vorticity fields at consecutive time frames, $k-1$, $k$, and $k+1$ as shown in Figure 1. It’s important to note that in this context, $k$ refers to the time frame within the entire dataset’s time span, rather than representing a specific time step during the forward or backward process in the diffusion model. This tri-channel approach enables the approximation of $\frac{\partial\omega}{\partial k}$ using finite difference methods.
@@
In the sampling process, the reverse Markov chain recovers the high-fidelity flow field starting from the low-fidelity observations $\tilde{\omega}$. This process inverts the forward diffusion, iteratively reconstructing the previous time step $\omega_{t-1}$ from a more noisy time step $\omega_{t}$, until the original flow field $\omega_{0}$ is retrieved:
$\omega_{t-1}=\frac{\eta_{t-1}}{\eta_{t}}\omega_{t}+\frac{\alpha_{t}}{\eta_{t}}f_{\theta}(\omega_{t},\tilde{\omega},t)+\kappa\sqrt{\frac{\eta_{t-1}}{\eta_{t}}\alpha_{t}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,I)$ (13)
We choose UNet as the underlying neural network architecture because of its exceptional performance in capturing complex spatial hierarchies inherent in image-like data, which is analogous to flow fields. With its encoding and decoding pathways, the architecture’s design is well-suited to preserving high-resolution details crucial for reconstructing the intricate structures found in vorticity fields, making it a proper choice for this task.